Key to the European species of Rhodocybe

(Basis: M.E. Noordeloos in Flora Agaricina Neerlandica (1988)) (Additions: M.E. Noordeloos in Karstenia 34, 43 (1994) (Section Rhodophana)) (Additional species: Rh. ardosiaca, peculiaris, stangliana, tillii)

1. Hymenial pseudocystidia present, very obvious, filled with yellowish, refringent strongly dextrinoid granules

2. Fr.b. pleurotoid . . . . . . . . . . . . . . Rhodocybe spec. (ND99016)

2. Fr.b. centrally stipitate

3. Cap concentrically cracking; smell and taste indistinct; sp. minutely nodulose-pustulate in outline . . . . . . . . . . . . . . . . . Rh. caelata 3. Cap smooth to minutely felty; smell and taste rancid-farinaceous; sp. with conspicuous hemispherical nodules . . . . . . . . Rh. ardosiaca

1. Hymenial pseudocystidia absent

4. Clamps present

5. Fr.b. tricholomatoid; stipe base with conspicuous, bulb-like mycelial mass . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Rh. stangliana

5. Fr.b. mycenoid to collybioid; base without bulb

6. Sp. (4−)4.5 − 7 × 3 − 4(−4.5) μm, Q̄ = 1.3 − 1.4; cap and stipe rather pale yellow brown . . . . . . . . . . . . . Rh. melleopallens

6. Sp. larger; cap and stipe darker

7. Cap and stipe honey-coloured to orange or reddish brown

8. Cap honey-coloured; sp. 6.5−8.5(−9)×(3.5−)4−5(−5.5) μm, Q = 1.5 − 1.9; sp. print greyish . . . . . . . Rh. griseospora 8. Cap vividly orange to red or red-brown; sp. 7−10×5−5.5 μm, Q̄ = 1.4 − 1.5; sp. print pinkish brown . . . . . Rh. nitellina

7. Cap and stipe dull coloured, brown, grey-brown or grey

9. Cap pale, greyish to whitish; sp. subglobose, (4.5−)5−6(−6.5)× (3.5−)4 − 4.5(−5) μm . . . . . . . . . . . . . . . Rh. peculiaris

9. Cap darker; sp. ellipsoid, larger

10. Cap and stipe dark grey or grey-brown; lam. with distinct grey tinge; sp. 7 − 10 × 4.5 − 6 μm …. Rh. hispanica 10. Cap and stipe dark brown; lam. sordidly pink when mature; sp. 6 − 8 × 4.5 − 5 μm . . . . . . . . . Rh. fuscofarinacea

4. Clamps absent

11. Fr.b. pleurotoid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Rh. tillii 11. Fr.b. centrally stipitate

12. Cap white, occasionally tinged cream colour at centre when old

Rh. fallax

12. Cap distinctly coloured

13. Cap flesh-coloured or reddish; stipe white or concolorous with cap . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Rh. gemina

13. Cap grey to grey-brown 14. Sp. 6−9×4−6 μm, ellipsoid to subglobose in side view, hardly

angular or pustulate; cheilocystidia present, 30−50×4−9 μm, usually septate . . . . . . . . . . . . . . . . . . . . . Rh. hirneola

14. Sp. 4.5 − 7(−8) μm long, distinctly pustulate; cheilocystidia absent

15. Smell none, taste mild . . . . . . . . . . . . . . . . Rh. parilis 15. Smell farinaceous, taste usually bitter . . Rh. popinalis

Version 1999/12/16